An airplane accelerates down a runway at 3.20 m/s^2 for 32.8s until is finally lifted off the ground. Determine the distance travelled before takeoff
1
Expert's answer
2018-09-27T12:08:09-0400
We can find the distance traveled before takeoff from the kinematic equation:
d=v_0 t+1/2 at^2,
here, d is the distance traveled before takeoff, v_0=0 is the initial velocity of the airplane, a is the acceleration of the airplane and t is the time.
Then, we get:
d=v_0 t+1/2 at^2=0+1/2∙3.20 m/s^2 ∙(32.8 s)^2=1721.3 m.